OpenStudy (anonymous):

Prove the Identity: 1+sinx-cos^2x/1+sinx=sinx

OpenStudy (anonymous):

Use the identity \(\sin^2x+\cos^2x=1\). Or, equivalently, \(\sin^2x=1-\cos^2x\). \[\begin{align*}\frac{1+\sin x-\cos^2x}{1+\sin x}&=\frac{1-\cos^2x+\sin x}{1+\sin x}\\ &=\frac{\sin^2x+\sin x}{1+\sin x} \end{align*}\] Then factor out a power of \(\sin x\) from the numerator: \[\begin{align*}\frac{1+\sin x-\cos^2x}{1+\sin x}&=\frac{1-\cos^2x+\sin x}{1+\sin x}\\ &=\frac{\sin^2x+\sin x}{1+\sin x}\\ &=\frac{\sin x(\sin x+1)}{1+\sin x} \end{align*}\] This is no different from what @Ahmad1 said the first time you asked this question, by the way.

OpenStudy (anonymous):

From here, \(\dfrac{\sin x+1}{1+\sin x}=\dfrac{1+\sin x}{1+\sin x}=1\), so you're left with just \(\sin x\).

OpenStudy (anonymous):

Wow........Muuuuuuuuch clearer. Thanks

OpenStudy (anonymous):

you're welcome